Home‎ > ‎Crew 164‎ > ‎

0304-DailySummaryReport

DAILY SUMMARY REPORT
SOL: 12
Person filling out Report: Mohammad IRANMANESH, XO
Summary Title: Goodbye Mars, you will be missed!

Mission Status: The overall operations mission performance is nominal. HAB has been cleaned and ready for crew rotation.

Sol Activity Summary: EVA#12. Repacking experiments and cleaning HAB. Inventory of food and supplies at the HAB. End of mission report. 

Look Ahead Plan: Ready for crew rotation.

Anomalies in work: /

Weather: cloudy in the morning but warm.

Crew Physical Status: No issue.

EVA: EVA in the morning to gather last set of data for projects that still needed measures (MOMA, Mind Wandering) and testing wide angle camera on the rover. All ATVs were refueled.

Reports to be filed: 
- Commander report
- Engineering report
- Journalist report 
- EVA #12 report

Support Requested: /